### v3

@@ C.3 Simulation: Heterogeneous coefficients
Table 3: Bias and standard error (s.e.) for the polynomial-regression FAT when the counterfactual is specified as in the left-most column. The time trend component is heterogeneous across individuals (top panel), with the addition of a cross-sectionally heterogeneous autoregressive component (bottom panel). Stationary initial condition for the AR($1$) component for each $i$.
C.4 Empirical replication: Overdose mortality and medical cannabis laws
[…]
@@ C.4 Empirical replication: Overdose mortality and medical cannabis laws
[…]
@@
[…]
We also compute placebo FAT by assuming that the law was adopted by each state either one year earlier or two years earlier, respectively, in the bottom panel of Figure 6. We consider these values because FAT in the top panel is computed using the outcome either one period before adoption (when $q=1$) or two periods before adoption (when $q=2$). Figure 6 shows the placebo FAT estimates computed with $R=q+1$, $q=1,2$. The forecast horizon is $h=1$, so the placebo FAT at lag 0 corresponds to FAT at forecast horizon 1 in the top panel. The figure shows that the estimated placebo FATs are statistically insignificant. Although this is not a test of our assumptions, these placebo estimates offer suggestive evidence that FAT estimates may be interpreted as the ATT across different forecast horizons.
@@
Figure 6: The effect of adopting legalized marijuana laws on overdose mortality rate. The top panel shows estimates of FAT as a function of forecast horizon (horizontal axis) and polynomial order (colored lines). FAT is computed with a variable number of pre-treatment time periods, $R=q+1,q=1,2$. The bottom panel shows estimates of placebo FAT as a function of lag (horizontal axis) and polynomial order (colored lines). Lag 0 is the actual adoption year. Placebo FAT is computed with a variable number of pre-treatment time periods, $R=q+1,q=1,2$.
C.5 Empirical replication: Refugees and far-right support
[…]
@@ C.5 Empirical replication: Refugees and far-right support
We include here placebo FAT estimates which we compute by assuming that the treatment took place in $2015$. The point here is to assess if the estimated placebo FAT is statistically not different from zero. We use a polynomial of degree $q=0$ when the pre-treatment year is $2013$ and a polynomial of degree $q=1$ when the pre-treatment years are $2012$ and $2013$. The placebo FATs are all statistically insignificant: the placebo FAT for the treated when $q=0$ is $-0.010$ with a standard error of $0.013$, while when $q=1$ it is $-0.009$ with a standard error of $0.029$. The placebo FAT for the control when $q=0$ is $-0.010$ with a standard error of $0.012$, while when $q=1$ it is $-0.008$ with a standard error of $0.017$. The corresponding placebo DFAT’s are then: 00 for $q=1$ with a standard error of $-0.001$ and $0.016$ for $q=1$ with a standard error of $0.034$.

### v4

@@ C.2 Simulation: Heterogeneous coefficients
Table 6: Bias and standard error (s.e.) for the Unobserved-Components FAT when the counterfactual is specified as in the left-most column. The time trend component is heterogeneous across individuals (top panel), with the addition of a cross-sectionally heterogeneous autoregressive component (bottom panel). Stationary initial condition for the AR(11) component for each ii.
C.3 Empirical replication: Overdose mortality and medical cannabis laws
We use data from Shover et al. (2019) to analyze the effect of adopting legalized medical cannabis laws on opioid overdose mortality in the US. Shover et al. (2019) contributes to the debate about whether the adoption of such laws has decreased overdose mortality (see, e.g., Bachhuber et al. 2014).
@@ C.3 Empirical replication: Overdose mortality and medical cannabis laws
Both Bachhuber et al. (2014) and Shover et al. (2019) use two-way fixed-effects estimators to analyze the effects of adopting legalized medical cannabis laws. In a staggered adoption setting, this estimation method produces biased results, e.g. Goodman-Bacon (2021). Therefore, we first redo the analysis to remove the bias of the original studies by using various methods, such as the staggered DiD approach of Callaway and Sant’Anna (2021) and the generalized synthetic control method of Xu (2017). Figure 3 uses the method of Callaway and Sant’Anna (2021) with not-yet-treated-states as control units, while Figure 4 uses the method of Xu (2017) to impute counterfactuals for each treated unit using a linear two-way fixed effects regression. This analysis finds an initial increase in overdose mortality and then a reversal, but neither effect is statistically significant.
@@
We then implement FAT. We start with a plot of the log mortality rate averaged across states as a function of time to adoption; see Figure 5, in order to get a sense for the time series properties of the outcome of interest. The outcome appears to be non-stationary over the pre-treatment period, so we choose the smallest possible estimation window $R=q+1$ to calculate FAT (as in (29)) and let the order of the polynomial be $q=1,2$. Figure 6 shows our FAT estimates for the ATT as a function of the forecast horizon. Our estimates are relatively stable across the polynomial orders, and our results show a slight increase in the overdose mortality rate. However, the increase does not appear to be statistically significant across the forecast horizon. Our results thus corroborate those of the approaches of Callaway and Sant’Anna (2021) and Xu (2017), but without using a control group.
We also compute placebo FAT by assuming that the law was adopted by each state either one year earlier or two years earlier, respectively, in Figure 6. We consider these values because FAT is computed using the outcome either one period before adoption (when $q=1$) or two periods before adoption (when $q=2$). Figure 6 shows the placebo FAT estimates computed with $R=q+1$, $q=1,2$. The forecast horizon is $h=1$. The figure shows that the estimated placebo FATs are statistically insignificant. Although this is not a test of our assumptions, these placebo estimates offer suggestive evidence that FAT estimates may be interpreted as the ATT across different forecast horizons.
@@
Figure 6: The effect of adopting legalized marijuana laws on overdose mortality rate. Estimates of FAT (to the right of the dashed vertical line) are shown as a function of forecast horizon (horizontal axis) and polynomial order (colored lines). FAT is computed with a variable number of pre-treatment time periods, $R=q+1,q=1,2$. Estimates of placebo FAT (to the left of the dashed vertical line) are shown as a function of lag (horizontal axis) and polynomial order (colored lines). Lag 0 is the actual adoption year. Placebo FAT is computed with a variable number of pre-treatment time periods, $R=q+1,q=1,2$.
C.4 Empirical replication: Refugees and far-right support
In this replication exercise, we use data from Dinas et al. (2019) which examines the relationship between refugee arrivals and support for the far right. Dinas et al. (2019) consider the case of Greece, and make use of the fact that some Greek islands (those close to the Turkish border) witnessed sudden and unexpected increases in the number of refugees during the summer of 2015, while other nearby Greek islands saw much more moderate inflows of refugees. The municipalities in the former Greek islands are considered treated, while the municipalities in the latter are considered controls. The authors use a standard DiD analysis to assess whether the treated municipalities were more supportive of the far-right Golden Dawn party in the September 2015 general election. The original data set contains a total of 96 municipalities, 48 of which were treated, and data on four elections: three pre-treatment elections in 2012, 2013, 2015, and one post-treatment election in 2016. The outcome of interest is the vote share for Golden Dawn (GD). Figure 7 shows the vote share for GD averaged across municipalities, treated and control, before and after the treatment time.
@@ C.4 Empirical replication: Refugees and far-right support
We include here placebo FAT estimates which we compute by assuming that the treatment took place in $2015$. The point here is to assess if the estimated placebo FAT is statistically indistinguishable from zero. We use a polynomial of degree $q=0$ when the pre-treatment year is $2013$ and a polynomial of degree $q=1$ when the pre-treatment years are $2012$ and $2013$. The placebo FATs are all statistically insignificant: the placebo FAT for the treated when $q=0$ is 0.010-0.010 with a standard error of $0.013$, while when $q=1$ it is 0.009-0.009 with a standard error of $0.029$. The placebo FAT for the control when $q=0$ is 0.010-0.010 with a standard error of $0.012$, while when $q=1$ it is 0.008-0.008 with a standard error of $0.017$. The corresponding placebo DFATs are then: 0 for $q=0$ with a standard error of $0.016$ and 0.001-0.001 for $q=1$ with a standard error of $0.034$.
@@
C.5 Additional simulation results
We consider the same DGP as that corresponding to Table 5 in Section 4.2. We report simulation results for a smaller sample size, $N=30$, in Table 8. This particular sample size is approximately that of two of our applications (in two of our applications, the sample size is $34$). The simulation results show that the large standard errors seen in the applications may be due to the small sample size. We note here that the size of the standard error of our estimator seen is comparable to that obtained in the studies that we replicate.
$R$
$q+1$ $q+2$ $q+3$ $q+4$ $q+5$
Stationary AR(1) $I_{1}=1,I_{2}=0=I_{3}$ $q=0$
bias -0.0001 -0.0023 0.0025 0.0011 0.0019
s.e. 0.2396 0.2139 0.2056 0.2056 0.2044
$q=1$
bias 0.0042 -0.0097 -0.0008 -0.0016
s.e. 0.4027 0.3105 0.2745 0.2574
$q=2$
bias 0.0252 -0.0144 -0.005
s.e. 0.7297 0.4874 0.3984
Stationary AR(1) + unit root $I_{1}=1=I_{2},I_{3}=0$ $q=0$
bias -0.0093 0.0003 0.0031 0.00045 -0.0015
s.e. 0.3102 0.3031 0.3118 0.3315 0.3467
$q=1$
bias -0.0285 -0.0149 -0.0005 -0.0023
s.e. 0.4966 0.4141 0.3641 0.3533
$q=2$
bias -0.0489 -0.045 -0.0156
s.e. 0.8543 0.6345 0.5282
Stationary AR(1) + linear trend $I_{1}=1=I_{3},I_{2}=0$ $q=0$
bias 0.9999 1.4977 2.0025 2.5011 3.0019
s.e. 0.2396 0.2139 0.2056 0.2056 0.2044
$q=1$
bias 0.0042 -0.0097 -0.0008 -0.0016
s.e. 0.4027 0.3105 0.2745 0.2574
$q=2$
bias 0.0252 -0.0144 -0.005
s.e. 0.7297 0.4874 0.3984
Stationary AR(1) + linear trend + unit root $I_{1}=I_{2}=I_{3}=1$ $q=0$
bias 0.9907 1.5003 2.0031 2.5004 3.0015
s.e. 0.3102 0.3031 0.3118 0.3315 0.3467
$q=1$
bias -0.0285 -0.0149 -0.0005 0.0023
s.e. 0.4966 0.4141 0.3641 0.3533
$q=2$
bias -0.0489 -0.045 -0.0156
s.e. 0.8543 0.6345 0.5282
Table 8: Bias and standard error (s.e.) for the Unobserved-Components FAT when the counterfactual is specified as indicated in the left-most column. Sample size $N=30$.